If it takes 6 hours to paint 5/7 of a truck, how long will it take to paint the whole truck? Thank you.
Answer by rfer(16322)

6/(5/7)=x/(7/7)
5/7x=7/7*6
5/7x=42/7
x=42/7*7/5
x=294/35
x=8.4 hrs